### v1

@@ Abstract
Polychoric correlation is often an important building block in the analysis of rating data, particularly for structural equation models. However, the commonly employed maximum likelihood (ML) estimator is highly susceptible to misspecification of the polychoric correlation model, for instance through violations of latent normality assumptions. We propose a novel estimator that is designed to be robust to partial misspecification of the polychoric model, that is, the model is only misspecified for an unknown fraction of observations, for instance (but not limited to) careless respondents. In contrast to existing literature, our estimator makes no assumption on the type or degree of model misspecification. It furthermore generalizes ML estimation and is consistent as well as asymptotically normally distributed. We demonstrate the robustness and practical usefulness of our estimator in simulation studies and an empirical application on a Big Five administration. In the latter, the polychoric correlation estimates of our estimator and ML differ substantially, which, after further inspection, is likely due to the presence of careless respondents that the estimator helps identify.
@@ 1 Introduction
To enhance accessibility and adoption by empirical researchers, an implementation of our proposed methodology in R (R Core Team,, 2024) is freely available in the package robcat (for “ROBust CATegorical data analysis”) at https://github.com/mwelz/robcat. [Replication files have been submitted for review together with this manuscript and will be made publicly available upon acceptance, with a link included here.]
@@ 4.2 Response probabilities under partial misspecification
It is worth addressing that there exist nonnormal distributions of the latent variables $(\xi,\eta)$ that, after discretization with the same thresholds, result in the same response probabilities as under latent normality (Foldnes & Grønneberg,, 2019). This implies that there may exist misspecifying distributions $H$ and (nonzero) misspecification fractions $\varepsilon$ under which the population probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are equal to the true population probabilities of the polychoric model, $p_{xy}(\bm{\theta}_{*})$, that is, $f_{\varepsilon}\left(x,y\right)=p_{xy}(\bm{\theta}_{*})$ for $\varepsilon>0$. In this situation, the polychoric model is misspecified, but the misspecification does not have consequences because the response probabilities remain unaffected so that there is no robustness issue with ML estimation. To avoid cumbersome notation in the theoretical analysis of our robust estimator, we assume consequential misspecification throughout this paper, that is, $f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*})$ whenever $\varepsilon>0$. However, it is silently understood that misspecification need not be consequential, in case of which there is no non-robustness issue and both the ML estimator and our robust estimator will be consistent for the true $\bm{\theta}_{*}$.
@@ 5.1 The estimator
The proposed estimator is a special case of a robust estimator for general models of categorical data introduced by Welz, (2024), and is based on the following idea. The empirical relative frequency of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$, denoted
@@
It follows that if the polychoric model is misspecified, there exists no parameter value $\bm{\theta}\in\bm{\Theta}$ whose associated model probability $p_{xy}(\bm{\theta})$ the nonparametric estimate $\widehat{f}_{N}(x,y)$ converges to. Hence, it is indicative of model misspecification if $\widehat{f}_{N}(x,y)$ does not converge to any polychoric model probability $p_{xy}(\bm{\theta})$, resulting in a discrepancy of $\widehat{f}_{N}(x,y)$ and $p_{xy}(\bm{\theta})$. This observation can be exploited in model fitting by minimizing the discrepancy between the empirical relative frequencies, $\widehat{f}_{N}(x,y)$, and theoretical model probabilities, $p_{xy}(\bm{\theta})$, to find the most accurate fit that can be achieved with the polychoric model for the observed data. Specifically, our estimator minimizes with respect to $\bm{\theta}$ the loss function
$L\left(\bm{\theta},\ \widehat{f}_{N}\right)=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\varphi\left(\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}\right)p_{xy}(\bm{\theta}),$ (9)
where $\varphi:[0,\infty)\to\mathbb{R}$ is a prespecified discrepancy function that will be defined momentarily. The proposed estimator $\widehat{\bm{\theta}}_{N}$ is given by the value minimizing the objective loss over parameter space $\bm{\Theta}$,
@@
For the discrepancy function choice $\varphi(z)=\varphi^{\mathrm{MLE}}(z)=z\log(z)$, it can be easily verified that $\widehat{\bm{\theta}}_{N}$ coincides with the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (4). In the following, we motivate a specific choice of discrepancy function $\varphi(\cdot)$ that makes the estimator $\widehat{\bm{\theta}}_{N}$ less susceptible to misspecification of the polychoric model while preserving equivalence with ML estimation in the absence of misspecification.
The centralized fraction between empirical relative frequencies and model probabilities, $\widehat{f}_{N}(x,y)/p_{xy}(\bm{\theta})-1$, is referred to as Pearson residual (Lindsay,, 1994) and can be interpreted as a goodness-of-fit measure. In our notation, it is simpler to renounce on subtracting the value 1 so that we call
$\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}$
@@
a decentralized Pearson residual (DPR), which takes values in $[0,+\infty)$. DPR values close to 1 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward 0 or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their DPR being away from 1, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
$\varphi(z)=\begin{cases}z\log(z)&\textnormal{ if }z\in[0,c],\\ z(\log(c)+1)&\textnormal{ if }z>c,\end{cases}$ (11)
@@
where $c\in[1,\infty]$ is a prespecified tuning constant that governs the estimator’s behavior at the DPR of each possible response. Figure 2 visualizes this function for the example choice $c=1.6$ as well as the ML discrepancy function function $\varphi^{\mathrm{MLE}}(z)=z\log(z)$.
@@
Figure 2: Visualization of the robust discrepancy function $\varphi(z)$ in (11) for $c=1.6$ (solid line) and the ML discrepancy function $\varphi^{\mathrm{MLE}}(z)=z\log(z)$ (dotted line).
@@
For the choice $c=+\infty$, minimizing the loss (9) is equivalent to maximizing the log-likelihood objective in (4), meaning that the estimator $\widehat{\bm{\theta}}_{N}$ is equal to $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for this choice of $c$. More specifically, if a DPR $z=\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}$ of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is such that $z\in[0,c]$ for fixed $c\geq 1$, then the estimation procedure behaves at this response like in classic ML estimation. As argued before, in the absence of misspecification, $\widehat{f}_{N}(x,y)$ converges to $p_{xy}(\bm{\theta}_{*})$ for all responses $(x,y)\in\mathcal{X}\times\mathcal{Y}$, therefore all DPRs are asymptotically equal to 1. Hence, if the polychoric model is correctly specified, then estimator $\widehat{\bm{\theta}}_{N}$ is asymptotically equivalent to the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for any tuning constant value $c\geq 1$. On the other hand, if a response’s DPR $z$ is larger than $c$, that is, $z>c\geq 1$, then the estimation procedure does not behave like in ML, but the response’s contribution to loss (9) is linear rather than super-linear like in ML (Figure 2). It follows that the influence of responses that cannot be fitted well by the polychoric model is downweighted to prevent them from dominating the fit. The tuning constant $c\geq 1$ is the “threshold” beyond which a DPR will be downweighted, so the choice thereof determines what is considered an insufficient fit. The closer to 1 the tuning constant $c$ is chosen, the more robust the estimator.
@@
With the proposed choice of $\varphi(\cdot)$, we stress that our estimator $\widehat{\bm{\theta}}_{N}$ in (10) has the same time complexity as ML, that is, $O\big{(}K_{X}\cdot K_{Y}\big{)}$, since one needs to calculate the Pearson residual of all $K_{X}\cdot K_{Y}$ possible responses for every candidate parameter value. Consequently, our proposed estimator does not incur any additional computational cost compared to ML.
@@ 5.2 Statistical properties
$\begin{split}&L\Big{(}\bm{\theta},f_{0}\Big{)}\\ &\quad=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\varphi\left(\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}\right)p_{xy}(\bm{\theta})\\ &\quad=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p_{xy}(\bm{\theta}_{*})\Bigg{[}{\log}\Bigg{(}\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}\Bigg{)}\mathds{1}\left\{\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}\in[0,c]\right\}+\left(\log(c)+1\right)\mathds{1}\left\{\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}>c\right\}\Bigg{]}\end{split}$
attains its global minimum of zero if and only if $\bm{\theta}=\bm{\theta}_{*}$, for any choice of $c\geq 1$. Thus, in the absence of misspecification, our estimator estimates the same quantity as the MLE, namely the true $\bm{\theta}_{*}$.
@@
However, in the presence of misspecification ($\varepsilon>0$), the estimand $\bm{\theta}_{0}$ is generally different from the true $\bm{\theta}_{*}$. The population estimand being different from the true value translates to biased estimates of the latter as a consequence of the misspecification. How much they are apart depends on the unknown fraction of misspecification $\varepsilon$, the unknown type of misspecification $H$, as well as the choice of tuning constant $c$ in $\varphi(\cdot)$. Mainly, the larger $\varepsilon$ (more severe misspecification) and $c$ (less downweighting of hard-to-fit responses), the further $\bm{\theta}_{0}$ is away from $\bm{\theta}_{*}$. Figure 3 illustrates this behavior for the polychoric correlation coefficient at an example misspecified distribution that is described in Section 4.1 and in which the true polychoric correlation under the correct model amounts to  $\rho_{*}=0.5$. For increasing misspecification fractions, the MLE ($c=+\infty)$ estimates a parameter value that is increasingly farther away from the true $\bm{\theta}_{*}$, where already a misspecification fraction of less than $\varepsilon=0.15$ suffices for a sign flip in the correlation coefficient. Conversely, choosing tuning constant $c$ to be near 1 results in a much less severe bias. For instance, even at misspecification degree $\varepsilon=0.3$, the difference between estimand and true value is less than approximately 0.1.
@@
It can be shown that under certain regularity conditions that do not restrict the degree or type of possible partial misspecification beyond $\varepsilon\in[0,0.5)$, the robust estimator $\widehat{\bm{\theta}}_{N}$ is consistent for estimand $\bm{\theta}_{0}$ as well as asymptotically normally distributed. With this limit theory, one can construct standard errors and confidence intervals for every element in $\bm{\theta}_{0}$. Importantly, in the absence of model misspecification (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix of the robust estimator is equal to that of the fully efficient ML estimator. Hence, if the model is correctly specified, then robust estimator and ML estimator are asymptotically first and second order equivalent. We refer to Appendix A for a rigorous exposition and discussion of the robust estimator’s asymptotic properties.
@@ 5.3 Implementation
An important user choice is that of the tuning constant $c$ in discrepancy function (11). The closer $c$ is to 1, the more robust the estimator will be to possible misspecification of the polychoric model (see, e.g., Figure 3). However, in the presence of model misspecification, the more robust the estimator is made, the larger its estimation variance becomes. This is an instance of a well-known tradeoff between robustness and efficiency (e.g., Huber & Ronchetti,, 2009). Here, the choice of $c$ governs the robustness-efficiency tradeoff. In simulation experiments, we found that the choice $c=1.6$ yielded a good compromise (see, e.g., Figure 3 for the robustness gain), so we choose to use this value for all applications in this paper. However, we acknowledge that a detailed study, possibly a theoretical one, is necessary to provide guidelines on the choice of $c$. We will explore this in future work.

### v2

@@ Abstract
Polychoric correlation is often an important building block in the analysis of rating data, particularly for structural equation models. However, the commonly employed maximum likelihood (ML) estimator is highly susceptible to misspecification of the polychoric correlation model, for instance through violations of latent normality assumptions. We propose a novel estimator that is designed to be robust to partial misspecification of the polychoric model, that is, the model is only misspecified for an unknown fraction of observations, for instance (but not limited to) careless respondents. In contrast to existing literature, our estimator makes no assumption on the type or degree of model misspecification. It furthermore generalizes ML estimation, is consistent as well as asymptotically normally distributed, and comes at no additional computational cost. We demonstrate the robustness and practical usefulness of our estimator in simulation studies and an empirical application on a Big Five administration. In the latter, the polychoric correlation estimates of our estimator and ML differ substantially, which, after further inspection, is likely due to the presence of careless respondents that the estimator helps identify.
@@ 1 Introduction
To enhance accessibility and adoption by empirical researchers, an implementation of our proposed methodology in R (R Core Team,, 2024) is freely available in the package robcat (for “ROBust CATegorical data analysis”) at https://github.com/mwelz/robcat.
@@ 4.2 Response probabilities under partial misspecification
It is worth addressing that there exist nonnormal distributions of the latent variables $(\xi,\eta)$ that, after discretization with the same thresholds, result in the same response probabilities as under latent normality (Foldnes & Grønneberg,, 2019). This implies that there may exist misspecifying distributions $H$ and (nonzero) misspecification fractions $\varepsilon$ under which the population probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are equal to the true population probabilities of the polychoric model, $p_{xy}(\bm{\theta}_{*})$, that is, $f_{\varepsilon}\left(x,y\right)=p_{xy}(\bm{\theta}_{*})$ for $\varepsilon>0$. In this situation, the polychoric model is misspecified, but the misspecification does not have consequences because the response probabilities remain unaffected so that there is no non-robustness issue with ML estimation. To avoid cumbersome notation in the theoretical analysis of our robust estimator, we assume consequential misspecification throughout this paper, that is, $f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*})$ whenever $\varepsilon>0$. However, it is silently understood that misspecification need not be consequential, in case of which there is no non-robustness issue and both the ML estimator and our robust estimator will be consistent for the true $\bm{\theta}_{*}$.
@@ 5.1 The estimator
The proposed estimator is a special case of a class of robust estimators for general models of categorical data called $C$-estimators (Welz,, 2024), and is based on the following idea. The empirical relative frequency of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$, denoted
@@
It follows that if the polychoric model is misspecified, there exists no parameter value $\bm{\theta}\in\bm{\Theta}$ whose associated model probability $p_{xy}(\bm{\theta})$ the nonparametric estimate $\widehat{f}_{N}(x,y)$ converges to. Hence, it is indicative of model misspecification if $\widehat{f}_{N}(x,y)$ does not converge to any polychoric model probability $p_{xy}(\bm{\theta})$, resulting in a discrepancy between $\widehat{f}_{N}(x,y)$ and $p_{xy}(\bm{\theta})$. This observation can be exploited in model fitting by minimizing the discrepancy between the empirical relative frequencies, $\widehat{f}_{N}(x,y)$, and theoretical model probabilities, $p_{xy}(\bm{\theta})$, to find the most accurate fit that can be achieved with the polychoric model for the observed data. Specifically, our estimator minimizes with respect to $\bm{\theta}$ the loss function
$L\left(\bm{\theta},\ \widehat{f}_{N}\right)=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\varphi\left(\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}-1\right)p_{xy}(\bm{\theta}),$ (9)
where $\varphi:[-1,\infty)\to\mathbb{R}$ is a prespecified discrepancy function that will be defined momentarily. The proposed estimator $\widehat{\bm{\theta}}_{N}$ is given by the value minimizing the objective loss over parameter space $\bm{\Theta}$,
@@
For the discrepancy function choice $\varphi(z)=\varphi^{\mathrm{MLE}}(z)=(z+1)\log(z+1)$, it can be easily verified that $\widehat{\bm{\theta}}_{N}$ coincides with the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ in (4). In the following, we motivate a specific choice of discrepancy function $\varphi(\cdot)$ that makes the estimator $\widehat{\bm{\theta}}_{N}$ less susceptible to misspecification of the polychoric model while preserving equivalence with ML estimation in the absence of misspecification.
The fraction between empirical relative frequencies and model probabilities with value 1 deducted,
$\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}-1,$
@@
is referred to as Pearson residual (PR, Lindsay,, 1994). It takes values in $[-1,+\infty)$ and can be interpreted as a goodness-of-fit measure. PR values close to 0 indicate a good fit between data and polychoric model at $\bm{\theta}$, whereas values toward $-1$ or +$\infty$+ ∞ indicate a poor fit because empirical response probabilities disagree with their model counterparts. To achieve robustness to misspecification of the polychoric model, responses that cannot be modeled well by the polychoric model, as indicated by their PR being away from 0, should receive less weight in the estimation procedure such that they do not over-proportionally affect the fit. Downweighting when necessary is achieved through a specific choice of discrepancy function $\varphi(\cdot)$ proposed by Welz, (2024), which is a special case of a function suggested by Ruckstuhl & Welsh, (2001). The discrepancy function reads
$\varphi(z)=\begin{cases}(z+1)\log(z+1)&\textnormal{ if }z\in[-1,c],\\ (z+1)(\log(c+1)+1)-c-1&\textnormal{ if }z>c,\end{cases}$ (11)
@@
where $c\in[0,\infty]$ is a prespecified tuning constant that governs the estimator’s behavior at the PR of each possible response. Figure 2 visualizes this function for the example choice $c=0.6$ as well as the ML discrepancy function function $\varphi^{\mathrm{MLE}}(z)=(z+1)\log(z+1)$. We stress that although the discrepancy function (11) can be negative, the loss function (9) is always nonnegative (Welz,, 2024).
@@
Figure 2: Visualization of the robust discrepancy function $\varphi(z)$ in (11) for $c=0.6$ (solid line) and the ML discrepancy function $\varphi^{\mathrm{MLE}}(z)=(z+1)\log(z+1)$ (dotted line).
@@
For the choice $c=+\infty$, minimizing the loss (9) is equivalent to maximizing the log-likelihood objective in (4), meaning that the estimator $\widehat{\bm{\theta}}_{N}$ is equal to $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for this choice of $c$. More specifically, if a Pearson residual $z=\frac{\widehat{f}_{N}(x,y)}{p_{xy}(\bm{\theta})}-1$ of a response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is such that $z\in[-1,c]$ for fixed $c\geq 0$, then the estimation procedure behaves at this response like in classic ML estimation. As argued before, in the absence of misspecification, $\widehat{f}_{N}(x,y)$ converges to $p_{xy}(\bm{\theta}_{*})$ for all responses $(x,y)\in\mathcal{X}\times\mathcal{Y}$, therefore all PRs are asymptotically equal to 0. Hence, if the polychoric model is correctly specified, then estimator $\widehat{\bm{\theta}}_{N}$ is asymptotically equivalent to the MLE $\widehat{\bm{\theta}}_{N}^{\mathrm{\ MLE}}$ for any tuning constant value $c\geq 0$. On the other hand, if a response’s PR $z$ is larger than $c$, that is, $z>c\geq 0$, then the estimation procedure does not behave like in ML, but the response’s contribution to loss (9) is linear rather than super-linear like in ML (Figure 2). It follows that the influence of responses that cannot be fitted well by the polychoric model is downweighted to prevent them from dominating the fit. The tuning constant $c\geq 0$ is the “threshold” beyond which a PR will be downweighted, so the choice thereof determines what is considered an insufficient fit. The closer to 0 the tuning constant $c$ is chosen, the more robust the estimator.
@@
With the proposed choice of $\varphi(\cdot)$, we stress that our estimator $\widehat{\bm{\theta}}_{N}$ in (10) has the same time complexity as ML, namely $O\big{(}K_{X}\cdot K_{Y}\big{)}$, since one needs to calculate the Pearson residual of all $K_{X}\cdot K_{Y}$ possible responses for every candidate parameter value. Consequently, our proposed estimator does not incur any additional computational cost compared to ML.
@@ 5.2 Statistical properties
$\begin{split}L\Big{(}\bm{\theta},f_{0}\Big{)}&=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\varphi\left(\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}-1\right)p_{xy}(\bm{\theta})\\ &=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p_{xy}(\bm{\theta}_{*}){\log}\Bigg{(}\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}\Bigg{)}\mathds{1}\left\{\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\}\\ &\quad+\Big{(}p_{xy}(\bm{\theta}_{*})\big{(}\log(c+1)+1\big{)}-p_{xy}(\bm{\theta})(c+1)\Big{)}\mathds{1}\left\{\frac{p_{xy}(\bm{\theta}_{*})}{p_{xy}(\bm{\theta})}-1>c\right\}\end{split}$
attains its global minimum of zero if and only if $\bm{\theta}=\bm{\theta}_{*}$, for any choice of $c\geq 0$. Thus, in the absence of misspecification, our estimator estimates the same quantity as the MLE, namely the true $\bm{\theta}_{*}$. We refer to Welz, (2024) for details.
@@
However, in the presence of misspecification ($\varepsilon>0$), the estimand $\bm{\theta}_{0}$ is generally different from the true $\bm{\theta}_{*}$. The population estimand being different from the true value translates to biased estimates of the latter as a consequence of the misspecification. How much they are apart depends on the unknown fraction of misspecification $\varepsilon$, the unknown type of misspecification $H$, as well as the choice of tuning constant $c$ in $\varphi(\cdot)$. Mainly, the larger $\varepsilon$ (more severe misspecification) and $c$ (less downweighting of hard-to-fit responses), the further $\bm{\theta}_{0}$ is away from $\bm{\theta}_{*}$. Figure 3 illustrates this behavior for the polychoric correlation coefficient at an example misspecified distribution that is described in Section 4.1 and in which the true polychoric correlation under the correct model amounts to  $\rho_{*}=0.5$. For increasing misspecification fractions, the MLE ($c=+\infty)$ estimates a parameter value that is increasingly farther away from the true $\bm{\theta}_{*}$, where already a misspecification fraction of less than $\varepsilon=0.15$ suffices for a sign flip in the correlation coefficient. Conversely, choosing tuning constant $c$ to be near 0 results in a much less severe bias. For instance, even at misspecification degree $\varepsilon=0.3$, the difference between estimand and true value is less than approximately 0.1.
@@
It can be shown that under certain regularity conditions that do not restrict the degree or type of possible partial misspecification beyond $\varepsilon\in[0,0.5)$, the robust estimator $\widehat{\bm{\theta}}_{N}$ is consistent for estimand $\bm{\theta}_{0}$ as well as asymptotically normally distributed. With this limit theory, one can construct standard errors and confidence intervals for every element in $\bm{\theta}_{0}$. Importantly, in the absence of model misspecification (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix of the robust estimator is equal to that of the fully efficient ML estimator as long as $c>0$. Hence, if the model is correctly specified, then robust estimator and ML estimator are asymptotically first and second order equivalent. We refer to Appendix A for a rigorous exposition and discussion of the robust estimator’s asymptotic properties.
@@ 5.3 Implementation
An important user choice is that of the tuning constant $c$ in discrepancy function (11). The closer $c$ is to 0, the more robust the estimator will be to possible misspecification of the polychoric model (see, e.g., Figure 3). On the other hand, in the presence of model misspecification, the more robust the estimator is made, the larger its estimation variance becomes. Moreover, if the model is correctly specified, then Welz, (2024) shows that the most robust choice, $c=0$, is associated with two drawbacks, namely asymptotic nonnormality as well as certain finite sample issues. To avoid these drawbacks, we suggest to choose a value slightly larger than 0. In simulation experiments, we found that the choice $c=0.6$ yields a good compromise between robustness and efficiency in case of misspecification while not suffering from finite sample issues if the model is correctly specified. We therefore chose to use this value for all applications in this paper. However, we acknowledge that a detailed study, possibly a theoretical one, is necessary to provide guidelines on the choice of $c$. We will explore this in future work.